OpenStudy (anonymous):

A model of a volcano has a scale of 2.5 inches to 1000.00 feet. If the model is 18 inches high, find the height of the volcano.

OpenStudy (texaschic101):

2.5 / 1000 = 18 / x cross multiply 2.5x = 18000 x = 18000/2.5 x = 7200 18 inches = 7200 ft

OpenStudy (anonymous):

Thank you